Required information (The following information applies to the questions displayed below.] Endless Mountain Company manufactuThe companys chief financial officer (CFO), in consultation with various managers across the organization has developed the11. The company uses a last-in, first-out (LIFO) inventory flow assumption. This means that the most recently purchased raw ma. Total fixed cost per year b. Total variable cost per unit sold C. Contribution margin per unit sold d. Unit sales to breakPrepare a budgeted variable costing income statement for 2017. Stop your computations at net operating income. Variable Costi
Answer #1

1.
a. Total Fixed Cost = $150000x4+(25000+64000+12000+8000+8000)x4 = $1068000

b. Total Variable cost per unit sold =3.5*3+0.25*18+0.25*3+1.25 = $17 per unit

c. Contribution margin per unit sold = Selling price - Variable cost per unit
= $32 - $17 = $15 per unit

d. Unit Sales to break even = Fixed Cost / Contribution margin per unit
= $1068000 / 15 = 71200 units in a year

Dollar Sales to break even = 71200 x $32 = $2278400

e. Margin of Safety = Actual Sales - Break even sales
= (12000+37000+15000+25000)*32 - 2278400 = $569600

f. Degree of Operating Leverage = Contribution Margin / Operating Income
= (12000+37000+15000+25000)*15 / (12000+37000+15000+25000)*15-1068000
= $1335000 / 267000 = 5

2.

Variable Costing Income Statement
Sales Revenue $         28,48,000
Less Variable Costs
Direct Material $   9,34,500 =89000*3.5*3
Direct labor $   4,00,500 =89000*0.25*18
Manufacturing Overhead $       66,750 =89000*0.25*3
Selling and administrative expenses $   1,11,250 =89000*1.25
Total Variable Costs $         15,13,000
Contribution Margin $         13,35,000
Less Fixed Expenses
Manufacturing Overhead $   6,00,000 =150000*4
Selling and administrative expenses $   4,68,000 =(25000+64000+12000+8000+8000)*4
Total Fixed Expenses $         10,68,000
Net Operating Income $            2,67,000